OpenStudy (yanasidlinskiy):

Here's another one that I need help with. This one I've been struggling the entire day and had no way to figure it out. Please help!. Thanx. Let f(x) = x^2 + 6 and g(x) =x+8/x. Find ( g o f)(-7).

OpenStudy (anonymous):

Yana banana...this is easy.

OpenStudy (yanasidlinskiy):

Here's what I have so far: \[(\frac{ f }{ g })(x)=\frac{ f(x) }{ g(x) }=\frac{ x ^{2}+6 }{ x+8 }/x\]

OpenStudy (yanasidlinskiy):

Not for me.

OpenStudy (anonymous):

So i think you start with the g function first, so put -7 in the g(x) function : -7 + 8/-7 = -57/7 and then put this value in f(x) function : (-57/7)^2 + 6 =72.3 (i think, I'm not really good at these either)

OpenStudy (anonymous):

Well..it LOOKED easy...I got confused...I need some sleep.

OpenStudy (yanasidlinskiy):

Ok. Go to sleep. I'm sure I can get help from someone hopefully.

OpenStudy (yanasidlinskiy):

-7 + 8/-7 = -57/7 \(\huge{\leftarrow}\) Where are you getting these numbers from?

OpenStudy (anonymous):

oh i was using a calculator

OpenStudy (yanasidlinskiy):

Nvm. Ok. I see where your getting it from. And 72.3 would be my answer? If it is, it's not in any of my answer choices.

Miracrown (miracrown):

g(x) is x + (8/x)?

OpenStudy (yanasidlinskiy):

G(x) = \[\frac{ x+8 }{ x }\]

OpenStudy (anonymous):

ohhhhhhhhh its for all of it

OpenStudy (anonymous):

i misread the equation

Miracrown (miracrown):

:)

OpenStudy (anonymous):

okay so im getting 1/57

Miracrown (miracrown):

 first we want to find g o f which is the same as g(f(x)) :